OpenStudy (anonymous):

given f(x)=-5+|x-3|,find f(1)-f(5) a. -14 b. -4 c. 2 d. 0 e. none of these

OpenStudy (anonymous):

ok when it says f(1) just integrate 1 where is

OpenStudy (anonymous):

1=x

OpenStudy (anonymous):

\[(-5+\left| ?-3 \right|)-(-5+\left| ?-3 \right|)\]

OpenStudy (anonymous):

once u calculate let me know

OpenStudy (anonymous):

\[(-5+\left| 1-3 \right|)-(-5+\left| 5-3 \right|)\]

OpenStudy (anonymous):

\[(-5+2)-(-5+2)\]

OpenStudy (anonymous):

so zero?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

thanks